# Synthesizing Privacy-Preserving Text Data via Finetuning without Finetuning Billion-Scale LLMs

### v1

@@ Abstract
Synthetic data offers a promising path to train models while preserving data privacy. Differentially private (DP) finetuning of large language models (LLMs) as data generator is effective, but is impractical when computation resources are limited. Meanwhile, prompt-based methods such as private evolution (Xie et al., 2024; Hou et al., 2024), depend heavily on the manual prompts, and ineffectively use private information in their iterative data selection process. To overcome these limitations, we propose CTCL (Data Synthesis with ConTrollability and CLustering), a novel framework for generating privacy-preserving synthetic data without extensive prompt engineering or billion-scale LLM finetuning. CTCL pretrains a lightweight 140M conditional generator and a clustering-based topic model on large-scale public data. To further adapt to the private domain, the generator is DP finetuned on private data for fine-grained textual information, while the topic model extracts a DP histogram representing distributional information. The DP generator then samples according to the DP histogram to synthesize a desired number of data examples. Evaluation across five diverse domains demonstrates the effectiveness of our framework, particularly in the strong privacy regime. Systematic ablation validates the design of each framework component and highlights the scalability of our approach.
@@ 1 Introduction
Many artificial intelligence (AI) applications improves their model performances by leveraging user data. For example, models are improved by adapting to the typing text in user’s mobile virtual keyboard (Hard et al., 2018; Xu et al., 2023), and aligning with user preference in a chatbot (OpenAI, 2024; Google, 2024; Llama Team, 2024). However, training models on user data raises privacy concerns, particularly in domains involving highly sensitive information, such as healthcare records (Milmo & Stacey, 2025) and chat messages (Hogan, 2025). Researchers have shown that the training data can be memorized and potentially extracted from models (Carlini et al., 2021; Nasr et al., 2023; Carlini et al., 2023). Synthesizing privacy-preserving user data has emerged as a promising approach to mitigating these privacy risks. A popular approach is to differentially-private (DP) finetune a generative language model (LM) on user data, followed by generating synthetic data using the finetuned model (Bommasani et al., 2019; Putta et al., 2022; Mattern et al., 2022a; Yue et al., 2023). Benefiting from the development of open-sourced billion-scale large language models (LLMs) such as Llama (Touvron et al., 2023), DP-finetuned generators have demonstrated effectiveness in the downstream classification (Kurakin et al., 2023) and instruction tuning tasks (Yu et al., 2024). However, DP finetuning is both computationally expensive and resource-intensive, because it requires per-sample gradient operations in every training batch and large batch size to get good privacy-utility trade-off (Ponomareva et al., 2023). This results in higher memory usage and slower training speeds compared to the non-DP finetuning. Moreover, when the user data are decentralized across their own devices, and no centralized data collection is allowed following the data minimization privacy principle (McMahan et al., 2017; Kairouz et al., 2021; Daly et al., 2024), the devices performing local computations typically lack the necessary resources to finetune billion-scale LLMs.
@@
We validate our framework across five diverse downstream domains, including the medical contexts and human-to-human conversations, covering both generative tasks evaluated by the next-word prediction accuracy, and the standard classification tasks. Our framework demonstrates significant advantages over previous approaches, particularly under strict privacy constraints. Through a comprehensive analysis, we highlight the importance of each component in our design, and demonstrate the scalability of CTCL compared to prompting-based methods such as PE (Xie et al., 2024).
@@ 2 Related Work
This line of research explores data synthesis using only LLM inference APIs, typically leveraging prompt engineering with domain-specific knowledge, such as specifying document structures or assuming roles (Wu et al., 2024a). The Private Evolution (PE) framework (Lin et al., 2024; Xie et al., 2024; Hou et al., 2024) integrates the private information into the synthetic data through an iterative sample selection process. Specifically, the API-generated outputs are selected based on their proximity to private data measured by the differentially private nearest neighbors (DP-NN). In this setup, DP-NN serves as the sole mechanism for extracting information from the private data, limiting the extent to which its information is fully captured. Furthermore, the synthetic data size (which determines the number of bins in the DP-NN histogram) is often constrained in order to better tolerate the DP noise (see discussion in Appendix B). For instance, the synthetic datasets in (Xie et al., 2024) contain typically 2,000 to 5,000 examples across the experiments. Unlike the prompting-based methods, our framework does not require prompt engineering and prior domain knowledge when applied to downstream data. Additionally, our synthetic dataset size is not constrained, offering significantly greater scalability compared to the PE approach. We discuss more related work including private inference in Appendix B.
@@ 3 CTCL Framework
In this work, we propose CTCL (Data Synthesis with ConTrollability and CLustering), a framework for generating synthetic private data without requiring billion-scale DP finetuning or domain-specific prompt engineering. Figure 1 gives an overview of CTCL.
@@ 3.1 CTCL-Generator
We perform continual pretraining on top of BART-base (Lewis, 2019), a 140M-parameter sequence-to-sequence LM previously pretrained in an unsupervised manner. The model’s encoder-decoder architecture is well suitable for conditioning on inputs through the encoder while generating outputs via the decoder. Optimization was performed using the AdamW optimizer with a batch size of 4096 and a cosine learning rate schedule starting at . The implementation of the pretraining is based on RedCoast (Tan et al., 2023) using bf16 mixed precision and the pretraining takes approximately 24 hours on 256 TPU-v4 cores (Jouppi et al., 2023).
@@ 3.2 CTCL-Topic
Specifically, we utilized the November 2023 version of Wikipedia, which contains over 6 million pages. A publicly available 20M-parameter document embedding model111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 was applied to the entire Wikipedia corpus, followed by HDBSCAN clustering (McInnes et al., 2017), resulting in the identification of 1,300 clusters, each treated as a distinct topic. To represent each topic, we employed the KeyBERT (Sharma & Li, 2019) to annotate 10 keywords for each cluster. The implementation of the pipeline above is based on BERTopic222https://maartengr.github.io/BERTopic/ (Grootendorst, 2022).
@@ 4.1.1 Downstream Tasks
Our experiments contain three generative tasks and two classification tasks. The downstream generative tasks are evaluated by the next-word prediction accuracy, which needs the synthetic data to preserve fine-grained textual information from the private data. In contrast, the downstream classification tasks usually rely on co-occurrence patterns between labels and words in the synthetic data. Therefore, generative tasks tend to be more challenging than classification tasks. 333Size of the training datasets can be found in Appendix C.
@@ 4.1.3 Hyperparameters
For all settings involving DP finetuning, we use DP-Adam for 2000 steps with a batch size of 4096, a gradient norm clip of 1.0, and a weight decay of 0.1. The learning rate follows a linear decay schedule with 100 warmup steps, and the peak learning rate is selected from the range based on validation performance. The privacy budget accounts for both DP model finetuning and the collection of DP topic histogram statistics. We apply a Gaussian noise multiplier of 10 to the DP topic histogram. The noise multipliers for DP finetuning vary across settings depending on the training data size and the presence of a topic histogram (see Appendix E). For the sample generation process, we generate 400K synthetic examples using nucleus sampling with top-p = 0.95 and a maximum sequence length of 512 tokens. For upper-bound experiments with -1.5B, we reduce the batch size to 256 to mitigate computational costs. The implementation of DP finetuning is based on RedCoast (Tan et al., 2023) using full fp32 precision.
@@ 4.2.1 Generative Tasks
Table 3 and Figure 2 present the results of three generative tasks444A complete result table is available in the Appendix D.. Our framework consistently outperforms baselines under different DP constraints and achieves performance close to the upper bounds. Moreover, as shown in Figure 2, the performance gap between our framework and the baselines widens under tighter privacy constraints (i.e., comparing the patterns of blue and yellow bars), highlighting its robustness. This can be attributed to our framework’s ability to simultaneously learn both high-level and fine-grained information from private data.
@@ 5 Conclusion
In this work, we propose a novel framework for synthesizing private domain data, which integrates a universal topic model with a lightweight 140M conditional language model. This framework captures both high-level, topic-specific information and fine-grained, context-sensitive details of the private domain in a modular and efficient manner. Through evaluations across five diverse downstream domains, we demonstrate that the synthesized data generated by our framework outperforms baseline methods, including vanilla differential privacy finetuning and prompting-based approaches such as private evolution.
Broader Impact
@@ Appendix B Supplementary Discussion on Related Work
Because DP finetuning can be expensive and sometimes impossible (e.g., for non-public models), this line of work explores data synthesis assuming only access to the LLM inference APIs. Simply relying on the high-level knowledge about a private domain to design proper LLM prompts is not enough to generate synthetic data that well represent the actual private domain (Wu et al., 2024a). The Private Evolution (PE) framework, initially developed by Lin et al. (2024) for the image domain, and later extended to the text domain in (Xie et al., 2024; Hou et al., 2024), proposes to “refine” (i.e., select good examples from) the current synthetic dataset according to the closeness of each example with respect to the private dataset. A similar idea is also explored by Zhao et al. (2024). The key idea behind PE is to measure closeness using DP nearest neighbor histogram: if an example is close to the private distribution, then it would receive a lot nearest neighbor votes from the private examples. PE starts with an initial dataset generated by the state-of-the-art LLMs via API access (using domain knowledge to design LLM prompts). PE then works by iteratively selecting good examples (measured by the DP nearest neighbor histogram) and using LLMs to generate more similar examples. In our experiments, we show that PE performs worse than DP finetuning especially for generative tasks, even when we are only allowed to finetune a small (million parameters) LMs.
@@
The size of the synthetic dataset in PE-based approaches is often constrained due to the sample-wise noise introduced by the differential-private nearest neighbor (DP-NN) process. In this setup, DP-NN serves as the only mechanism for incorporating the information from the private data into the synthetic dataset. Specifically, DP-NN identifies the nearest neighbors of each private data sample within the synthetic dataset. To preserve privacy, DP noise is added to each synthetic sample. In this process, each synthetic sample acts as a bin, with its count indicating how many private data samples identify it as their nearest neighbor. Noise is then applied to these counts to ensure privacy. However, this process requires each bin to contain a sufficiently large number of counts; otherwise, the noise overwhelms the signal, making it difficult to distinguish between zero and small counts. Consequently, the synthetic dataset size must be limited, as an excessive number of bins would make DP-NN ineffective. To overcome this limitation, a variation process is often employed, where LLM APIs are prompted to generate additional samples based on the initially selected subset. However, since the private data does not directly influence this generation process, the final synthetic dataset theoretically contains no more information from the private data than the small subset originally selected.
@@
Another line of work explores generating private-preserving few-shot examples for in-context learning (ICL), e.g., (Tang et al., 2024; Wu et al., 2024b; Duan et al., 2023). Given a reasonable DP guarantee, these methods can only generate a limited amount of synthetic data, e.g., a few for ICL, or at most thousand (Amin et al., 2024). By contrast, our method (based on DP finetuning) can generate a much larger dataset of synthetic examples.
@@ Appendix C Dataset Sizes
Appendix D Full Results of Generative Tasks

### v2

@@ Abstract
Synthetic data offers a promising path to train models while preserving data privacy. Differentially private (DP) finetuning of large language models (LLMs) as data generator is effective, but is impractical when computation resources are limited. Meanwhile, prompt-based methods such as private evolution (Xie et al., 2024; Hou et al., 2024) depend heavily on the manual prompts, and ineffectively use private information in their iterative data selection process. To overcome these limitations, we propose CTCL (Data Synthesis with ConTrollability and CLustering), a novel framework for generating privacy-preserving synthetic data without extensive prompt engineering or billion-scale LLM finetuning. CTCL pretrains a lightweight 140M conditional generator and a clustering-based topic model on large-scale public data. To further adapt to the private domain, the generator is DP finetuned on private data for fine-grained textual information, while the topic model extracts a DP histogram representing distributional information. The DP generator then samples according to the DP histogram to synthesize a desired number of data examples. Evaluation across five diverse domains demonstrates the effectiveness of our framework, particularly in the strong privacy regime. Systematic ablation validates the design of each framework component and highlights the scalability of our approach.
@@ 1 Introduction
Many artificial intelligence (AI) applications improve their model performances by leveraging user data. For example, models are improved by adapting to the typing text in user’s mobile virtual keyboard (Hard et al., 2018; Xu et al., 2023), and aligning with user preference in a chatbot (OpenAI, 2024; Google, 2024; Llama Team, 2024). However, training models on user data raises privacy concerns, particularly in domains involving highly sensitive information, such as healthcare records (Milmo & Stacey, 2025) and chat messages (Silva, 2025). Researchers have shown that the training data can be memorized and potentially extracted from models (Carlini et al., 2021; Nasr et al., 2023; Carlini et al., 2023). Synthesizing privacy-preserving user data has emerged as a promising approach to mitigating these privacy risks. A popular approach is to differentially-private (DP) finetune a generative language model (LM) on user data, followed by generating synthetic data using the finetuned model (Bommasani et al., 2019; Putta et al., 2022; Mattern et al., 2022a; Yue et al., 2023). Benefiting from the development of open-sourced billion-scale large language models (LLMs) such as Llama (Touvron et al., 2023), DP-finetuned generators have demonstrated effectiveness in the downstream classification (Kurakin et al., 2023) and instruction tuning tasks (Yu et al., 2024). However, DP finetuning is both computationally expensive and resource-intensive, because it requires per-sample gradient operations in every training batch and large batch size to get good privacy-utility trade-off (Ponomareva et al., 2023). This results in higher memory usage and slower training speeds compared to the non-DP finetuning. Moreover, when the user data are decentralized across their own devices, and no centralized data collection is allowed following the data minimization privacy principle (McMahan et al., 2017; Kairouz et al., 2021; Daly et al., 2024), the devices performing local computations typically lack the necessary resources to finetune billion-scale LLMs.
@@
We validate our framework across five diverse downstream domains, including the medical contexts and human-to-human conversations, covering both generative tasks evaluated by the next-word prediction accuracy, and the standard classification tasks. Our framework demonstrates significant advantages over previous approaches, particularly under strict privacy constraints. Through a comprehensive analysis, we highlight the importance of each component in our design, and demonstrate the scalability of CTCL compared to prompting-based methods such as PE (Xie et al., 2024). 111Code available at https://github.com/tanyuqian/synthetic-private-data
@@ 2 Related Work
This line of research explores data synthesis using only LLM inference APIs, typically leveraging prompt engineering with domain-specific knowledge, such as specifying document structures or assuming roles (Wu et al., 2024a; Zhang et al., 2025). The Private Evolution (PE) framework (Lin et al., 2024; Xie et al., 2024; Hou et al., 2024) integrates the private information into the synthetic data through an iterative sample selection process. Specifically, the API-generated outputs are selected based on their proximity to private data measured by the differentially private nearest neighbors (DP-NN). In this setup, DP-NN serves as the sole mechanism for extracting information from the private data, limiting the extent to which its information is fully captured. Furthermore, the synthetic data size (which determines the number of bins in the DP-NN histogram) is often constrained in order to better tolerate the DP noise (see discussion in Appendix B). For instance, the synthetic datasets in (Xie et al., 2024) contain typically 2,000 to 5,000 examples across the experiments. Unlike the prompting-based methods, our framework does not require prompt engineering and prior domain knowledge when applied to downstream data. Additionally, our synthetic dataset size is not constrained, offering significantly greater scalability compared to the PE approach. We discuss more related work including private inference in Appendix B.
@@ 3 CTCL Framework
In this work, we propose CTCL (Data Synthesis with ConTrollability and CLustering), a framework for generating synthetic private data without requiring billion-scale DP finetuning or domain-specific prompt engineering. Figure 1 and Appendix H give an overview of CTCL.
@@ 3.1 CTCL-Generator
We perform continual pretraining on top of BART-base (Lewis, 2019), a 140M-parameter sequence-to-sequence LM previously pretrained in an unsupervised manner. The model’s encoder-decoder architecture is well suitable for conditioning on inputs through the encoder while generating outputs via the decoder. Optimization was performed using the AdamW optimizer with a batch size of 4096 and a cosine learning rate schedule starting at . The implementation of the pretraining is based on RedCoast (Tan et al., 2024) using bf16 mixed precision and the pretraining takes approximately 24 hours on 256 TPU-v4 cores (Jouppi et al., 2023).
@@ 3.2 CTCL-Topic
Specifically, we utilized the November 2023 version of Wikipedia, which contains over 6 million pages. A publicly available 20M-parameter document embedding model222https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 was applied to the entire Wikipedia corpus, followed by HDBSCAN clustering (McInnes et al., 2017), resulting in the identification of 1,300 clusters, each treated as a distinct topic. To represent each topic, we employed the KeyBERT (Sharma & Li, 2019) to annotate 10 keywords for each cluster. The implementation of the pipeline above is based on BERTopic333https://maartengr.github.io/BERTopic/ (Grootendorst, 2022).
@@ 4.1.1 Downstream Tasks
Our experiments contain three generative tasks and two classification tasks444Training data sizes can be found in Appendix C.. The downstream generative tasks are evaluated by the next-word prediction accuracy, which needs the synthetic data to preserve fine-grained textual information from the private data. In contrast, the downstream classification tasks usually rely on co-occurrence patterns between labels and words in the synthetic data. Therefore, generative tasks tend to be more challenging than classification tasks.
@@ 4.1.3 Hyperparameters
For all settings involving DP finetuning, we use DP-Adam for 2000 steps with a batch size of 4096, a gradient norm clip of 1.0, and a weight decay of 0.1. The learning rate follows a linear decay schedule with 100 warmup steps, and the peak learning rate is selected from the range based on validation performance. The privacy budget accounts for both DP model finetuning and the collection of DP topic histogram statistics. We apply a Gaussian noise multiplier of 10 to the DP topic histogram. The noise multipliers for DP finetuning vary across settings depending on the training data size and the presence of a topic histogram 555More details of noise multipliers and privacy budget allocation can be found in Appendix E and D.. For the sample generation process, we generate 400K synthetic examples using nucleus sampling with top-p = 0.95 and a maximum sequence length of 512 tokens. For upper-bound experiments with -1.5B, we reduce the batch size to 256 to mitigate computational costs. The implementation of DP finetuning is based on RedCoast (Tan et al., 2024) using full fp32 precision.
@@ 4.2.1 Generative Tasks
Table 3 and Figure 2 present the results of three generative tasks 666A complete result table is available in Appendix K.. Our framework consistently outperforms baselines under different DP constraints and achieves performance close to the upper bounds. Moreover, as shown in Figure 2, the performance gap between our framework and the baselines widens under tighter privacy constraints (i.e., comparing the patterns of blue and yellow bars), highlighting its robustness. This can be attributed to our framework’s ability to simultaneously learn both high-level and fine-grained information from private data.
@@ 5 Conclusion
In this work, we propose a novel framework CTCL for synthesizing private domain data, which integrates a universal topic model with a lightweight 140M conditional language model. This framework captures both high-level, topic-specific information and fine-grained, context-sensitive details of the private domain in a modular and efficient manner. Through evaluations across five diverse downstream domains, we demonstrate that the synthesized data generated by CTCL outperforms baseline methods, including vanilla differential privacy finetuning and prompting-based approaches such as private evolution.
While the proposed CTCL framework outperforms existing privacy-preserving methods in the downstream tasks, one limitation is that the synthetic data generated by our 140M-parameter still lack the fluency and coherence compared to those generated by prompting the latest billion-size LLMs. How to leverage LLMs for data synthesis in the resource-constrained setting (the setting which this paper focuses on), would be one important area for future exploration.
Another interesting direction for future work is to generalize our framework to the multi-modal settings, e.g., data synthesis for the vision-language tasks. This requires appropriate architectural adaptations, e.g., by replacing language models with diffusion-based models for image generation.
Impact Statement
@@ Appendix B Supplementary Discussion on Related Work
Because DP finetuning can be expensive and sometimes impossible (e.g., for non-public models), this line of work explores data synthesis assuming only access to the LLM inference APIs. Simply relying on the high-level knowledge about a private domain to design proper LLM prompts is not enough to generate synthetic data that well represent the actual private domain (Wu et al., 2024a). Using a small DP language model to filter the initial synthetic data makes the data more similar to private data, but does not close the distribution gap for training a model for the downstream task in the private domain (Wu et al., 2024a; Zhang et al., 2025).
The Private Evolution (PE) framework, initially developed by Lin et al. (2024) for the image domain, and later extended to the text domain in (Xie et al., 2024; Hou et al., 2024), proposes to “refine” (i.e., select good examples from) the current synthetic dataset according to the closeness of each example with respect to the private dataset. A similar idea is also explored by Zhao et al. (2024). The key idea behind PE is to measure closeness using DP nearest neighbor histogram: if an example is close to the private distribution, then it would receive a lot nearest neighbor votes from the private examples. PE starts with an initial dataset generated by the state-of-the-art LLMs via API access (using domain knowledge to design LLM prompts). PE then works by iteratively selecting good examples (measured by the DP nearest neighbor histogram) and using LLMs to generate more similar examples. In our experiments, we show that PE performs worse than DP finetuning especially for generative tasks, even when we are only allowed to finetune a small (million parameters) LMs.
@@
The size of the synthetic dataset in PE-based approaches is often constrained due to the sample-wise noise introduced by the differential-private nearest neighbor (DP-NN) process. In this setup, DP-NN serves as the only mechanism for incorporating the information from the private data into the synthetic dataset. Specifically, DP-NN identifies the nearest neighbors of each private data sample within the synthetic dataset. To preserve privacy, DP noise is added to each synthetic sample. In this process, each synthetic sample acts as a bin, with its count indicating how many private data samples identify it as their nearest neighbor. Noise is then applied to these counts to ensure privacy. However, this process requires each bin to contain a sufficiently large number of counts; otherwise, the noise overwhelms the signal, making it difficult to distinguish between zero and small counts. Consequently, the synthetic dataset size must be limited, as an excessive number of bins would make DP-NN ineffective. To overcome this limitation, a variation process is often employed, where LLM APIs are prompted to generate additional samples based on the initially selected subset. However, since the private data does not directly influence this generation process, the final synthetic dataset theoretically contains no more information from the private data than the small subset originally selected. A few concurrent work (Hou et al., 2025; Zou et al., 2025) further improved the PE algorithm; while we did not directly compare to their experimental results, a lot of our discussion can still apply to these improved PE methods.
@@
Another line of work explores generating private-preserving few-shot examples for in-context learning (ICL), e.g., (Tang et al., 2024; Wu et al., 2024b; Duan et al., 2023). Given a reasonable DP guarantee, these methods can only generate a limited amount of synthetic data, e.g., a few for ICL, or a few thousands (Amin et al., 2024, 2025). By contrast, our method (based on DP finetuning) can generate a much larger dataset of synthetic examples.
@@
Conditional generation.
In addition to privacy-preserving methods for synthesizing data for a specific domain, conditional generation is also explored to improve the diversity and generative quality of models in less privacy sensitive applications. We now discuss a few recent work in large language models. In concurrent work, Gao et al. (2025) discussed metadata conditioning for pre-trainig large language models; DeSalvo et al. (2025) developed soft prompt as condition to improve synthetic data generation.
@@ Appendix C Dataset Sizes
Appendix D Privacy Budget Allocation
Our privacy budget allocation follows the “Post-Generation Resample” algorithm proposed by Yu et al. (2024), which also has the DP-histogram and DP-finetune components. For the DP-histogram step, we adopt the same configuration—adding a Gaussian noise on every histogram bin with a standard deviation of 10. The overall privacy budget () is computed as the composition of both steps using the standard DP accounting library777https://github.com/google/differential-privacy/tree/main/python/dp_accounting/dp_accounting/pld.
To illustrate, we provide the individual values in the table below for the DP-histogram and DP-finetuning steps on the PubMed dataset (75K training samples), under composed privacy budgets of = 4, 2, and 1, and . As shown below, the DP-histogram step consistently accounts for a small portion of the overall budget—for instance, for DP-histogram versus for DP-finetuning when the composed is 4. This trend aligns with the observations in (Yu et al., 2024) and highlights that our CTCL-Topic design achieves strong performance while consuming only a small portion of the privacy budget.
| (Composed) | (DP-Histogram) | (DP-Finetune) |
| 4 | 0.39 | 3.96 |
| 2 | 0.39 | 1.94 |
| 1 | 0.39 | 0.9 |
Additionally, we find that changing the privacy allocation to the DP-histogram step only has very small impact to the DP-finetuning step. For example, in the = 4 setting on PubMed, increasing the Gaussian noise in the histogram step from 10 to 20 alters the DP-finetuning marginally—from 3.96 to 3.99. This corresponds to a small change in the noise multiplier for DP-finetuning (from 3.03 to 3.02).
Appendix E Noise Multipliers
The table below gives the noise multipliers used when DP finetuning LMs in our experiments. Following (Yue et al., 2023), we set , where is the size of private training set. Given a desired -DP guarantee, the noise multipliers are computed using the standard dp_accounting package (DP Team, 2022). As pointed out in Appendix A, we use DP-Adam for DP finetuning and follow the standard Gaussian mechanism to obtain -DP guarantee. Compared to the “Vanilla DP Finetune” approach, the noise multiplier used by our method is slightly larger, because we need to allocate a small portion of the privacy budget to the DP topic histogram (see Appendix D). Besides, -1.5B has much smaller noise multipliers because we reduce the training batch size from 4096 to 256 to save computational resources. For other non-DP training hyperparameters, see §4.1.3.
| PubMed | ||||
| Vanilla DP Finetune ( and ) | 0 | 3.01 | 5.49 | 10.3 |
| -1.5B (reduced batch size) | 0 | 0.63 | 0.77 | 0.97 |
| Ours | 0 | 3.03 | 5.63 | 11.33 |
| Chatbot Arena | ||||
| Vanilla DP Finetune ( and ) | 0 | 1.47 | 2.5 | 4.58 |
| -1.5B (reduced batch size) | 0 | 0.56 | 0.67 | 0.78 |
| Ours | 0 | 1.48 | 2.57 | 5.08 |
| Multi-Session Chat | ||||
| Vanilla DP Finetune ( and ) | 0 | 11.38 | 21.01 | 39.41 |
| -1.5B (reduced batch size) | 0 | 1.00 | 1.52 | 2.61 |
| Ours | 0 | 11.45 | 21.47 | 42.70 |
| Yelp | ||||
| Vanilla DP Finetune ( and ) | 0 | 0.63 | 0.77 | 0.91 |
| -1.5B (reduced batch size) | 0 | 0.51 | 0.6 | 0.69 |
| Ours | 0 | 0.63 | 0.77 | 0.94 |
| OpenReview | ||||
| Vanilla DP Finetune ( and ) | 0 | 23.3 | 42.87 | 80.05 |
| -1.5B (reduced batch size) | 0 | 1.64 | 2.81 | 5.11 |
| Ours | 0 | 23.44 | 43.72 | 86.05 |
Appendix F Topic Model’s Coverage
We choose Wikipedia as the basis for our topic model because it is large-scale, high-quality, and semantically diverse. Most real-world documents with meaningful content would fall within Wikipedia’s domain. In practice, it is rare to encounter texts that are entirely outside of its scope. To illustrate the coverage of our topic model, the table below presents topics sampled from five diverse datasets used in our experiments. These examples demonstrate that the model captures a broad range of universal topics:
| Dataset | Sample topic 1 | Sample Topic 2 |
| PubMed | microscopy/electron/… | rna/mir/gene/… |
| Chatbot Arena | gameplay/rpg/wii/… | eruptions/volcano/… |
| Multi-Session Chat | comedian/presenter/… | oscar/nominations/… |
| Yelp | restaurant/diner/… | theater/cinemas/… |
| OpenReview | grammar/syntax/… | computational/turing/… |
The “out-of-domain” text is also considered and processed in our experiments. Specifically, in our topic model, there is an “unclassified” bin for documents that are not sufficiently close to any learned topic. The table below reports the number of unclassified samples in each dataset, along with example such texts:
| Dataset | Training size | Unclassified samples | Unclassified sample |
| PubMed | 75,316 | 0 | |
| Chatbot Arena | 180,000 | 86 | hello. i have no idea who am i, maybe you can help? |
| Multi-Session Chat | 17,940 | 0 | |
| Yelp | 1,939,290 | 4 | I am unable to provide check number. |
| OpenReview | 8,396 | 0 |
As shown in Table 11, the number of unclassified samples is neglectable across all datasets. Furthermore, those that do fall into the unclassified bin typically contain little to no substantive content, confirming the broad coverage of our topic model.
Appendix G Computational Cost and Tradeoffs
Our pretraining pipeline involves three main components:
Data Curation: Running inference using a 2B LLM on a 430M pretraining documents.
Generator Pretraining: Training a 140M conditional LM on 430M (condition, document) pairs.
Topic Modeling: Generating embeddings for 6M Wikipedia documents with a 20M document embedding model, followed by a HDBSCAN clustering.
Among these, data curation for pretraining dominates the computational cost due to the scale of LLM inference. This remains lighter than or comparable to the cost of the pretraining of a 2B LLM. Note that after we release our pretrained CTCL-Generator and CTCL-Topic, other uses can skip the pretraining step (i.e., do not need to pay the pretraining costs), and directly finetune the model on their private data.
For the synthetic data generation stage, our approach samples from a small 140M model after finetuning on the private data. By contrast, existing PE-based methods often involve significant API costs that are proportionally to the sample size. For instance, in the experiments of AUG-PE (Xie et al., 2024), they synthesize 2,000 samples using evolution iterations, each with variations, totaling ChatGPT requests.
Appendix H Pseudo-code
CTCL-Topic Development
CTCL-Generator Pretraining
Private Topic Histogram Construction
DP Finetuning CTCL-Generator
Appendix I Discrepancy in Constructed Conditions between Pretraining and Finetuning
In our pretraining stage (§3.1), the condition is extracted for each document from Gemma 2B, while in the learning of private data, the condition is from the topic model. The distinction between pretraining and finetuning conditions is intentional and serves two main purposes: enhancing model generalizability and adhering to privacy constraints. During the pretraining stage, using the 2B LLM to extract aspects is for increasing the diversity and flexibility of the aspects the model can handle, enabling stronger controllability at inference time. Otherwise, “keywords” becomes the only aspect in the pretraining.
On the other hand, not also using Gemma 2B in the finetuning time is for privacy consideration. Specifically, if Gemma 2B were applied directly to private data, the extracted conditions would be in free-text format. This raises our concerns about sending the private information to the synthetic data generation process. In contrast, our framework ensures differential privacy by using a noised, topic-wise histogram instead. This approach preserves key high-level topic information of the private domain while satisfying privacy requirements.
To demonstrate the value of the flexible conditioning during pretraining, consider a scenario where one has a bit prior knowledge about the downstream domain (e.g., that the documents are dialogues). Our design enables such weak supervision to be incorporated, providing a better initialization for subsequent DP-finetuning. Below are examples of model generations before any finetuning, controlled only by document type, showcasing the better initialization:
| Condition 1 |
| Document Type: News |
| Keywords: Surfing, World Championship, Young Athletes |
| Generation |
| The 15-year-old, who is a member of the World Surfing Association (WSA) and the World Surf League (WSL), won the World Championships for the first time in his career. […] |
| Condition 2 |
| Document Type: Chat |
| Keywords: Surfing, World Championship, Young Athletes |
| Generation |
| How do you feel about the current state of surfing in the U.S. right now? |
| There are a lot of great young athletes out there right now, and it’s going to be a really exciting year for them. […] |
| Condition 3 |
| Document Type: Blog Post |
| Keywords: Surfing, World Championship, Young Athletes |
| Generation |
| I’ve been fortunate enough to have the opportunity to meet some of the most talented young athletes in the world, and I’m […] |
Appendix J Qualitative Analysis
We provide synthetic data examples in addition to Table 6.
| BART-Base (Downstream Model Performance: 30.5%): |
| Sample 1: We explored the relationship between molecular interaction, NCT-2 and NCT-3 (NCT-1), NCT-4, NCT and NCT–1.5 (NCT), NCT-, NCT-6, and NCT[4]. In a recent clinical trial, we described an enzyme in NCT-10 that enabled novel processes to explore novel approaches for NCT- 2.3 to NCT-III. NCT was applied to NCT and the NCT-8 was used to explore the implications of NCT on NCT in NCT. NCT-11 was used to describe NCT-7 to NCT preparation for NCT preparation. NCT and PCT were compared with NCT-16 with NCT and […] |
| Sample 2: The authors of this paper present a unique method for understanding the mechanism of the effects of the mechanism. It aims to investigate whether the mechanism used by the methods involved in the study was used to determine the mechanism of treatment, including the mechanism of its regulation. The analysis was conducted on an open-source basis and did not rely on the traditional methods of the primary investigation, such as the original design for the method used, which looked at the mechanisms used to […] |
| -1.5B (Upper Bound, Downstream Model Performance: 37.7%) |
| Sample 1: The ability of leptin to induce weight loss, to stimulate ectothermic thermogenesis, and to augment activity of the AMPK system and the AMPK-dependent lipoprotein lipase activity, was examined. Circulating concentrations of leptin were assessed in the femoral adipose fat pad of the lean and obese rats. Serum leptin levels and AMPK mRNA expression were assessed in sera obtained from lean and obese female rats fed diets enriched with fat or carbohydrate. The leptin binding to the leptin receptor […] |
| Sample 2: Nuclear stress hormone (NTH) levels increased in patients with type 2 diabetes mellitus (T2D) following therapy with metformin [13] or insulin (e.g., lispro-Lipid, insulin-C3G or insulin-G3F) [14, 15]. NTH was not increased in patients enrolled to receive metformin or lispro (or in patients taking insulin-C-based drugs (i.e., lisulin and galantangulin, alemtuhusin) [16], suggesting the insulin treatment itself might cause patients to become insulin sensitive (i.g., via altered insulin action). […] |
| AUG-PE + GPT-3.5 (Downstream Model Performance: 30.3%) |
| Sample 1: Deep learning, in combination with artificial intelligence, has emerged as a powerful tool in the field of medical imaging. This study explores the application of deep learning algorithms in the identification of various medical conditions. By leveraging neural networks, this approach provides unparalleled accuracy across different modalities, enhancing our understanding of pathological processes. The integration of deep learning in medical imaging has the potential to revolutionize diagnostic processes […] |
| Sample 2: Investigating synaptic plasticity in the context of neurological networks investigation, a cohort of patients with neurodegenerative disorders was closely monitored. The efficiency of synaptic transmission was evaluated, considering the underlying mechanisms particularly focusing on the composition of scaffolding proteins. Insights into the impact of adipocyte signaling on synaptic function were uncovered. The research further characterized the dynamics of synaptic plasticity and illuminated how […] |
| Ours (Downstream Model Performance: 35.9%): |
| Sample 1: To develop a therapeutic formula to reduce rates of morbidity that occur in people with a combination of cardiovascular problems. We used a multi-state, multidisciplinary approach to the research of the clinical manifestations of cardiovascular problems with the introduction of a biocontrol. Using a multi-choice, multi-pronged approach we evaluated a combination of the three major elements of cardiovascular problems: chronic diseases; vascular disease that occurs in chronic conditions in chronic […] |
| Sample 2: This study describes the mechanism of action of different mechanisms to be performed to evaluate the efficacy and quality of control activities at three different levels. A cross-hatch method was used to investigate the efficacy and health outcomes of different mechanisms in all three levels of intervention. The outcome was compared with the outcomes obtained by a crosshatch method. A total of 20 control methods used during this study included 7 standard I/A for four trials. All intervention levels […] |
Appendix K Full Results of Generative Tasks